Hello, please could you help me with a question.
I need to find the surface area of a cylinder when the radius is 6cm and the length is 21 cm,
Please could you show the full step by step guide as i need to show the working out.
Many thanks

Answer by nycsub_teacher(90)

You need a formula and it is
surface area = 2(πr^2) + 2(πr)(h), where r = radius, h = height
and π = pi = 3.14
I guess by length you meant to type height.
Let SA = surface area of cynlinder for short.
SA = 2(3.14)(6)^2 + 2(3.14)(21)
SA = 226.08 + 131.88
SA = 357.96cm^2